# A Survey on Quantum Machine Learning: Basics, Current Trends, Challenges, Opportunities, and the Road Ahead

### v3

@@ Abstract
Furthermore, we explore practical applications of QML, illustrating how it can be leveraged to solve real-world problems more efficiently than classical ML methods. This paper serves as a valuable resource for readers seeking to understand the current state-of-the-art techniques in the QML field, offering a solid foundation to embark on further exploration and development in this rapidly evolving area.
@@ 1 Introduction
The high computational demands of modern ML models necessitate advanced hardware development. According to Moore’s law, the number of transistors on an integrated circuit doubles approximately every two years. However, this trend has reached its limits with current high-end CPUs and GPUs (Gustafson, 2011). This physical saturation restricts computational power, leading to delays in processing, developmental, and scientific discovery within the ML community. For instance, training Large Language Models with hundreds of billions of parameters and trillions of tokens is extremely compute-intensive (OpenAI, 2023). Such tasks require massive investments in time and hardware resources, which only a few high-end companies can afford. To overcome these physical limits and support further discoveries, there is an urgent need to explore new technological avenues of hardware systems that can enhance the computational efficiency for solving given real-world problems by closely simulating and comprehending them.
@@
Our work 2025 54 191
One of the most promising solutions to this bottleneck is Quantum Computing (QC). Initially proposed by Feynmann (1982) and further developed by Preskill (2018; 2023), QC systems exploit quantum mechanical phenomena to significantly improve performance and information processing compared to classical systems. Unlike classical systems that struggle to encompass natural processes, quantum computers operate on similar principles to those found in nature. This similarity suggests that QC could help us better understand natural phenomena. Moreover, QC has the potential to solve problems more cost-efficiently through optimizations, time savings, and environmentally friendly designs.
The Quantum Machine Learning (QML) paradigm represents an excellent opportunity for researchers and industries to achieve remarkable discoveries and design efficient solutions for complex real-world problems. QML systems, driven towards practicality and improved performance over classical systems, open new avenues for the community to discover, build, and align their designs across different levels of the quantum stack (Schuld & Killoran, 2022). This integration of QC and ML could lead to groundbreaking advancements and a deeper understanding of the world around us.
@@ 2.1 Understanding the Qubit
Understanding the qubit as an abstract mathematical object is crucial for the development of quantum circuits. Although the physical realization of qubits is important, the fundamental properties and specifications of qubits are dictated by quantum mechanical principles and remain consistent regardless of their physical form. Unlike a classical bit, which can only be in one of two states (0 or 1) at any given time, a qubit can exist in a superposition of both 0 and 1 simultaneously. This means that a qubit can exist in a superposition of both 0 and 1 states simultaneously (see Figure 1b). This quantum mechanical property of representing any qubit state as a linear combination of basis states, called superposition, allows a qubit to embody a range of probabilities for being in either state, reflecting its inherent quantum nature. This abstract representation of a qubit is consistent across any physical realization of the qubit, whether it is implemented using superconducting circuits, trapped ions, or any other technology.
@@
The Bloch sphere is an intuitive visual representation of a two-level (two states per qubit) single qubit system. As shown in Figure 2, it represents the pure and mixed states of a qubit. The pure states are distributed over the surface of the sphere, and all the internal states within the sphere are referred to as mixed-qubit states. It has three axes with their respective basis states on opposite ends.
@@ 2.2 Superposition
Before delving into the details of the superposition principle, we discuss two fundamental concepts of quantum mechanics, the Wave–Particle duality and the Max-Born rule.
The superposition principle is one of the most intriguing and counterintuitive aspects of quantum mechanics. It states that a quantum system can exist in multiple states simultaneously until it is measured. The space in which these states exist is called Hilbert space, a mathematical framework that allows the description of quantum states as vectors. A common physical interpretation of the superposition principle involves the spin of a particle, which can be in a superposition of “spin up” and “spin down” states or any combination thereof. Similarly, the energy levels of free electrons in an atom illustrate superposition. An electron can exist in a superposition of the ground state (|$\ket{0}$) and an excited state (|$\ket{1}$), with possible intermediate states.
@@ 2.2.2 Two-Qubit Superposition
On the contrary, a quantum system for the same problem will result in the same solution for identifying the position (state) of the object, but the approach is different. Based on quantum mechanics principles, the quantum system identifies the solution that results in a probabilistic distribution of the possible solutions to the given problem, i.e., in this case, the probabilistic positions of the object’s placement. The quantum system’s solution is that the object is at position B (state $01$). However, it is only $60\%$ confident that it is there and considers other possible outcomes. The mechanisms that quantum systems follow to understand and find the solution are much closer to the real world than classical systems.
@@ 2.3 Quantum Gates
Quantum logic gates are one of the essential parts of a quantum computer and are the building blocks of all quantum algorithms. These gates are mathematically described by unitary matrices, and their action is always logically reversible. The gates correspond to operators used to manipulate the quantum state of qubits. The only condition on a valid gate (and its corresponding quantum operator) is that it should be unitary. By unitary we mean that it meets the $U\textsuperscript{\textdagger}=U^{-1}$ condition, where $U$ is the gate operation matrix and $U\textsuperscript{\textdagger}$ is the adjoint of $U$. This property also ensures that it is a reversible gate operation (Barenco et al., 1995).
@@ 2.4 Quantum Circuits
A quantum circuit is a collection of quantum gates interconnected by quantum wires. The actual structure of a quantum circuit, the number and the types of gates, as well as the interconnection scheme, are dictated by the unitary transformation, U, executed by the circuit (Marinescu & Marinescu, 2012b). A circuit is an independent module composed of gates, wires, and qubits arranged in a certain way to perform a task. Hence, its functionality is similar to what we refer to as an algorithm. A collection of circuits are independent functional quantum units using the output of a circuit as the input of another one to form a complete algorithm for the solution of the given problem (Muñoz Coreas & Thapliyal, 2022) (see Figure 7).
@@ 2.5 Quantum Correlations
In contrast, quantum mechanics introduces a probabilistic nature to correlation. Observables in a quantum system do not have a single, definite outcome (Adesso et al., 2016). Instead, they have a range of possible values, known as eigenvalues, which are associated with certain probabilities. These eigenvalues are the result of the system being represented by a Hermitian matrix. Thus, even without any perturbation, measurements of a quantum system can yield different results over time. These varying outcomes are described by eigenfunctions that exist within a Hilbert space, a mathematical framework that encompasses all possible states of the system.
@@ 2.5.1 Quantum Entanglement
Quantum entanglement is a central concept in quantum information theory and represents one of the most counterintuitive aspects of quantum mechanics. Entanglement occurs when two particles become linked such that the state of one particle instantaneously influences the state of the other, regardless of the distance separating them (Bussandri & Lamberti, 2021; Cacciapuoti et al., 2020; Illiano et al., 2022). This phenomenon, which Albert Einstein famously referred to as a “spooky action at a distance”, defies the classical notion of local realism (Einstein et al., 1971).
@@ 2.5.2 Quantum Decoherence
Quantum decoherence is the process by which a quantum system loses its coherence due to interactions with the external environment. This loss of coherence manifests as the system’s transition from a pure quantum state to a mixed state, effectively causing a collapse of the wave function (Schlosshauer, 2019). Decoherence can be viewed as the leakage of information from the quantum system into the environment, or as environmental interference disrupting the quantum state.
@@ 2.6 Quantum Noise
Leading from the idea of coherence and decoherence, quantum noise makes the implementation of large-scale and reliable quantum computers with low error rates extremely challenging. Therefore, it is imperative to properly characterize the noise in quantum systems and devise mitigation techniques to detect and correct the errors (Shaib et al., 2021).
2.6.1 Uncertainty Principle
Quantum noise can be partially understood through the lens of Heisenberg’s uncertainty principle. This principle asserts that certain pairs of physical properties, such as position and momentum, cannot both be known to arbitrary precision simultaneously. The more precisely one property is known, the less precisely the other can be determined. This inherent uncertainty in quantum mechanics leads to quantum noise, which manifests as fluctuations in the quantum state of a system.
$\Delta x\cdot\Delta p\geq\frac{\hbar}{2}$ (2)
@@ 2.6.3 Error Correction
Due to the no-cloning theorem (Wootters & Zurek, 1982), creating a perfect copy of a quantum state is impossible. Hence, Quantum Error Correction (QEC) codes introduce redundancy by spreading the information of a single qubit onto an entangled state of multiple qubits. With this approach, it is possible to perform multi-qubit (syndrome) measurements to extract the information about the error without altering the quantum information. Typical QEC schemes are the three-qubit bit flip code (Peres, 1985), the three-qubit phase flip code (Nielsen & Chuang, 2010a) and Shor’s nine qubit code (Shor, 1995).
@@ 2.6.4 Error Mitigation
Despite sharing the same goal as QEC to reduce the impact of quantum noise, Quantum Error Mitigation (QEM) techniques operate differently. While QEC aims to restore the correct value after the error occurs, QEM aims to reduce or suppress the errors that occur during computation. Moreover, as shown in Figure 11, QEC completely removes the noise, while QEM keeps a small amount of noise under control. A scenario where QEC is applied, called ‘hard regime’, is theoretically ideal but difficult to realize. The other scenario where QEM is employed, called ‘easy regime’, is experimentally easier to implement but at the price of lower qubit protection. For QEM deployment, Mitiq (LaRose et al., 2022) is an open-source error mitigation package in Python that implements QEM techniques like zero-noise extrapolation (Giurgica-Tiron et al., 2020) and probabilistic error cancellation (van den Berg et al., 2023) for quantum computers.
@@ 2.6.5 Fault-Tolerant Quantum Computing
Fault tolerance is a property that enables a system to continue operating not only in normal conditions but also in the presence of hardware or software failures. While there is not often a direct comparison between classical and quantum information processing, the principles of fault tolerance in classical systems are easily transferable.
@@
2.7 Quantum Stack
Quantum computers operate under a fundamentally different set of rules than classical computers and can solve certain problems faster by exploiting quantum effects. The development stack in Figure 12 shows an abstraction of the quantum stages formulating a generalized quantum stack. Starting from the Hardware or qubit level, the abstractions go up to the application level composed of Quantum Algorithms. Since QC is at an early development stage, industries primarily focus on the Quantum Hardware level, where the reliability of measurements plays a crucial role in designing and choosing the physical mediums. While simulated quantum hardware can deal with perfect qubits, realistic quantum hardware can be implemented with superconducting qubits, ion trapped qubits, quantum dots, and neutral atom qubits (Bertels et al., 2021).
@@ 3 Benefits of Quantum Computing
This section presents an overview of the benefits of QC over classical systems. It provides the necessary motivations to conduct further research and development of quantum computers.
3.1 Quantum Dimensionality Reduction
A common way to reveal properties of an unknown quantum state, having many copies of a system in that state, is to measure different observables and to analyze the results statistically. The unknown quantum state can play an active role in this analysis. Given multiple copies of a quantum system with its respective density matrix, it is possible to perform the unitary transformation. Its result generates quantum coherence among different copies of the system. Popular techniques for dimensionality reduction are the Principal Component Analysis (PCA) (Lloyd et al., 2014), where the eigenvectors corresponding to the unknown state’s eigenvalues state are revealed in a very fast manner, and the Quantum Linear Discriminant Analysis Dimensionality Reduction (QLDADR) (Yu et al., 2023), where the vectors having a high-dimensional feature space are projected onto a low-dimensional feature space.
3.2 Quantum Supremacy for Solving Complex Problems

### v4

@@ Abstract
Furthermore, we explore practical applications of QML, illustrating how it can be leveraged to solve real-world problems more efficiently than classical ML methods. This survey aims to consolidate the current landscape of QML and outline key opportunities and challenges for future research.
@@ 1 Introduction
The high computational demands of modern ML models necessitate advanced hardware development. According to Moore’s law, the number of transistors on an integrated circuit doubles approximately every two years (Gustafson & Padua, 2011). However, recent analyses indicate that while Moore’s Law has historically driven exponential growth in transistor density, this trend has decelerated in recent years, prompting discussions about its continued applicability in the face of physical and economic challenges (Zhang, 2022). This physical saturation restricts computational power, leading to delays in processing, developmental, and scientific discovery within the ML community. For instance, training Large Language Models with hundreds of billions of parameters and trillions of tokens is extremely compute-intensive (OpenAI, 2023). Such tasks require massive investments in time and hardware resources, which only a few high-end companies can afford. To overcome these physical limits and support further discoveries, there is an urgent need to explore new technological avenues of hardware systems that can enhance the computational efficiency for solving given real-world problems by closely simulating and comprehending them.
@@
Our work 2025 61 235
One of the most promising solutions to this bottleneck is Quantum Computing (QC). Initially proposed by Feynmann (1982) and further developed by Preskill (2018; 2023), QC systems exploit quantum mechanical phenomena to significantly improve performance and information processing compared to classical systems. Unlike classical systems that struggle to capture the complexity of natural processes, quantum computers operate on similar principles to those found in nature. This similarity suggests that QC could help us better understand natural phenomena. Moreover, QC has the potential to solve problems more cost-efficiently by enabling advanced optimizations and reducing computation time.
The Quantum Machine Learning (QML) paradigm represents an excellent opportunity for researchers and industries to achieve remarkable discoveries and design efficient solutions for complex real-world problems. QML extends classical ML by leveraging quantum effects, such as superposition and entanglement, to process data in exponentially large feature spaces (Lloyd et al., 2013; Biamonte et al., 2017; Schuld et al., 2014). QML systems, driven towards practicality and improved performance over classical systems, open new avenues for the community to discover, build, and align their designs across different levels of the quantum stack (Schuld & Killoran, 2022). This integration of QC and ML could lead to groundbreaking advancements and a deeper understanding of the world around us.
@@ 2.1 Understanding the Qubit
Understanding the qubit as an abstract mathematical object is crucial for the development of quantum circuits. Although the physical realization of qubits is important, the fundamental properties and specifications of qubits are dictated by quantum mechanical principles and remain consistent regardless of their physical form. Unlike a classical bit, which can only be in one of two states (0 or 1) at any given time, a qubit can exist in a superposition of both 0 and 1 simultaneously (see Figure 1b).
Mathematically, a qubit is represented as a unit vector in a two-dimensional complex Hilbert space (see Equation 1).
|$\ket{\psi}=\alpha\ket{0}+\beta\ket{1},\quad\text{where }\alpha,\beta\in\mathbb{C},\quad|\alpha|^{2}+|\beta|^{2}=1$ (1)
Here, |$\ket{0}$ and |$\ket{1}$ are the computational basis states, and the coefficients $\alpha$ and $\beta$ are complex probability amplitudes. The condition $|\alpha|^{2}+|\beta|^{2}=1$ ensures normalization, according to the Max-Born rule.
This quantum mechanical property of representing any qubit state as a linear combination of basis states, called superposition, allows a qubit to embody a range of probabilities for being in either state, reflecting its inherent quantum nature. This abstract representation of a qubit is consistent across any physical realization of the qubit, whether it is implemented using superconducting circuits, trapped ions, or any other technology.
@@
As shown in Figure 2, the Bloch sphere provides a geometric representation of a qubit’s state, where any pure qubit state corresponds to a point on the surface of the sphere. This visualization aids in understanding qubit manipulations and quantum gate operations.
@@ 2.2 Superposition
The superposition principle is one of the most intriguing and counterintuitive aspects of quantum mechanics. It states that a quantum system can exist in multiple states simultaneously until it is measured. The space in which these states exist is called Hilbert space, a mathematical framework that allows the description of quantum states as vectors. The principle of superposition allows a qubit to exist simultaneously in multiple states until measured. Upon measurement, the qubit collapses to one of the basis states, |$\ket{0}$ or |$\ket{1}$, with probabilities determined by the squared magnitudes of the respective amplitudes. For example, a qubit in the state |$\ket{\psi}=\frac{1}{\sqrt{2}}(\ket{0}+\ket{1})$ has equal probabilities of being measured in either basis state.
@@ 2.2.2 Two-Qubit Superposition
On the contrary, a quantum system solving the same problem yields a probabilistic distribution over possible measurement outcomes due to the intrinsic uncertainty of quantum states, rather than simply classical randomness. This uncertainty originates from the principles of quantum mechanics, particularly the concept of superposition, and differs fundamentally from the statistical uncertainty described by classical thermodynamics or classical probability theory. When measured, the quantum state collapses into one of its eigenstates, such as position B (state |$\ket{01}$), with a certain probability (e.g., 60%). Unlike classical uncertainty, which arises from incomplete knowledge of a deterministic state, quantum uncertainty is inherent to the quantum state itself and persists even with complete knowledge of the system’s wavefunction.
@@ 2.3 Quantum Gates
Quantum logic gates are one of the essential parts of a quantum computer and are the building blocks of all quantum algorithms. Quantum gates are mathematically described by unitary operators that manipulate qubit states. By unitary we mean that it meets the $U\textsuperscript{\textdagger}=U^{-1}$ condition, where $U$ is the gate operation matrix and $U\textsuperscript{\textdagger}$ is the adjoint of $U$. This property also ensures that it is a reversible gate operation (Barenco et al., 1995). Common single-qubit gates include the Pauli-X, Pauli-Y, Pauli-Z, Rotation-X, Rotation-Y, Rotation-Z, and Hadamard gates. Multi-qubit gates, such as the Controlled-NOT (CNOT) gate, enable entanglement between qubits.
@@ 2.4 Quantum Circuits
Quantum circuits are sequences of quantum gates applied to qubits to perform computations. An example of a simple quantum circuit is shown in Figure 7, illustrating the application of gates and measurement operations. The actual structure of a quantum circuit, the number and the types of gates, as well as the interconnection scheme, are dictated by the unitary transformation, U, executed by the circuit (Marinescu & Marinescu, 2012b). A circuit is an independent module composed of gates arranged in a certain way across qubits to perform a task. Hence, its functionality is similar to what we refer to as an algorithm. A collection of circuits are independent functional quantum units using the output of a circuit as the input of another one to form a complete algorithm for the solution of the given problem (Muñoz Coreas & Thapliyal, 2022).
@@ 2.5 Quantum Correlations
In contrast, quantum mechanics introduces a probabilistic nature to correlation. Observables in a quantum system do not have a single, definite outcome (Adesso et al., 2016). Instead, they are represented by mathematical operators whose possible measurement outcomes are described by their eigenvalues (Griffiths & Schroeter, 2018). Each observable corresponds to a Hermitian (self-adjoint) operator, ensuring these eigenvalues are real numbers and physically measurable (Sakurai & Napolitano, 2020). When measuring a quantum state, the system collapses probabilistically into one of the eigenstates associated with these eigenvalues. This probabilistic behavior arises not merely from randomness, as seen in classical thermodynamics, but fundamentally from the quantum principle of superposition, where a quantum state can exist as a linear combination of multiple eigenstates prior to measurement (Ballentine, 2014). Consequently, repeated measurements of identically prepared quantum systems generally yield varying results according to well-defined probability distributions inherent to their quantum state representation in a Hilbert space (Griffiths & Schroeter, 2018).
@@ 2.5.1 Quantum Entanglement
Quantum entanglement is a central concept in quantum information theory and represents one of the most counterintuitive aspects of quantum mechanics. Entanglement occurs when two particles become linked such that the state of one particle instantaneously influences the state of the other, regardless of the distance separating them (Bussandri & Lamberti, 2021; Cacciapuoti et al., 2020; Illiano et al., 2022). Formally, the entanglement is a quantum phenomenon where the states of two or more qubits become correlated such that the state of one qubit cannot be described independently of the others. Measurement of one qubit in an entangled pair instantaneously determines the state of the other, regardless of the distance separating them. This phenomenon, which Albert Einstein famously referred to as a “spooky action at a distance”, defies the classical notion of local realism (Einstein et al., 1971).
@@ 2.5.2 Quantum Decoherence
Quantum decoherence is the process by which a quantum system loses its coherence due to interactions with the external environment, leading to errors in quantum computations. This loss of coherence manifests as the system’s transition from a pure quantum state to a mixed state, effectively causing a collapse of the wave function (Schlosshauer, 2019). Decoherence can be viewed as the leakage of information from the quantum system into the environment, or as environmental interference disrupting the quantum state.
@@ 2.6 Quantum Noise
Leading from the idea of coherence and decoherence, quantum noise encompasses various error sources (with corresponding noise channels), including bit-flip, phase-flip, amplitude damping, phase damping, and depolarizing errors. Quantum noise in realistic devices arises when qubits interact with uncontrolled environments, leading to non-unitary evolution that cannot be described by simple Schrödinger dynamics (Nielsen & Chuang, 2010b). Due to noise, the implementation of large-scale and reliable quantum computers with low error rates becomes extremely challenging. Therefore, it is imperative to properly characterize the noise in quantum systems and devise mitigation techniques to detect and correct the errors (Shaib et al., 2021).
In contrast to pure-state descriptions, density matrices capture classical and quantum uncertainties, encoding both mixedness and coherence. Noise channels map density matrices to more mixed states (higher entropy), degrading off-diagonal elements that represent quantum coherence (Nielsen & Chuang, 2010b; Lidar & Birgitta Whaley, 2003). This loss of coherence limits circuit depth and algorithmic performance on NISQ devices, making it essential to quantify noise strength via measures such as diamond norms or average gate fidelities (Magesan et al., 2011).
@@ 2.6.1 Noisy Intermediate-Scale Quantum Era
Contemporary quantum hardware remains limited in scale and fidelity. For example, IBM’s Eagle processor contains 127 superconducting qubits, and a recent IBM chip (Osprey) achieved 433 qubits (IBM, ), while IonQ’s latest Forte trapped-ion system operates with 32 qubits (with coherence times exceeding 1 second) (Chen et al., 2023). However, superconducting qubit devices typically have coherence times on the order of only tens of microseconds and two-qubit gate error rates around $10^{-3}$$10^{-2}$ per gate, whereas trapped-ion qubits can maintain quantum coherence for durations on the order of seconds with single- and two-qubit gate errors on the order of $10^{-4}$$10^{-3}$ (Chen et al., 2023). Furthermore, superconducting architectures generally allow only limited connectivity between qubits (each qubit usually interacts directly with only 2–3 neighbors in a heavy-hex lattice topology), which means that additional swap operations are required to entangle distant qubits. Each such operation adds noise and overhead. These hardware constraints, such as limited qubit count, finite coherence, gate errors, and restricted connectivity, impose strict limits on the depth and size of quantum circuits that can be executed reliably. As a consequence, quantum algorithms in the NISQ era must be shallow and noise-aware, and even then their performance and scalability are significantly curtailed by hardware noise and errors. For instance, a state-of-the-art trapped-ion system can execute on the order of 500 two-qubit gate operations in a single circuit before decoherence or errors become prohibitive, whereas many superconducting processors can only perform on the order of tens of two-qubit gates before accumulated errors overwhelm the results. Such comparisons illustrate the gap between current hardware capabilities and the requirements for practically useful quantum algorithms. . While not yet capable of fault-tolerant quantum computation, NISQ devices enable the exploration of quantum algorithms and applications in areas such as optimization, chemistry, and machine learning.
@@ 2.6.2 Error Correction
Due to the no-cloning theorem (Wootters & Zurek, 1982), creating a perfect copy of a quantum state is impossible. Hence, Quantum Error Correction (QEC) codes introduce redundancy by spreading the information of a single qubit onto an entangled state of multiple physical qubits. With this approach, it is possible to perform multi-qubit (syndrome) measurements to extract the information about the error without altering the quantum information. Typical QEC schemes are the three-qubit bit flip code (Peres, 1985), the three-qubit phase flip code (Nielsen & Chuang, 2010a) and Shor’s nine qubit code (Shor, 1995).
@@ 2.6.3 Error Mitigation
Despite sharing the same goal as QEC to reduce the impact of quantum noise, Quantum Error Mitigation (QEM) techniques operate differently. While QEC aims to restore the correct value after the error occurs, QEM aims to reduce or suppress the errors that occur during computation, without full error correction. Moreover, as shown in Figure 11, QEC completely removes the noise, while QEM keeps a small amount of noise under control. A scenario where QEC is applied, called ‘hard regime’, is theoretically ideal but difficult to realize. The other scenario where QEM is employed, called ‘easy regime’, is experimentally easier to implement but at the price of lower qubit protection. For QEM deployment, Mitiq (LaRose et al., 2022) is an open-source error mitigation package in Python that implements QEM techniques like zero-noise extrapolation (Giurgica-Tiron et al., 2020) and probabilistic error cancellation (van den Berg et al., 2023) for quantum computers.
@@ 2.6.4 Fault-Tolerant Quantum Computing
Fault tolerance is a property that enables a system to continue operating not only in normal conditions but also in the presence of hardware or software failures. Advancements in hardware and error correction are essential for transitioning from the NISQ era to scalable, fault-tolerant quantum computing. While there is not often a direct comparison between classical and quantum information processing, the principles of fault tolerance in classical systems are easily transferable.
@@ 2.7 Quantum Stack
To facilitate understanding quantum computing systems, it is helpful to consider them as structured in a hierarchical stack of abstraction layers, similar to classical computing architectures. As depicted in Figure 12, the quantum computing stack is generally composed of several distinct levels: quantum hardware (physical qubits and their implementation), quantum control (gate operations and measurement), quantum compilation (translation of quantum algorithms into hardware-specific gate sequences), and quantum algorithms and applications (high-level problem-solving procedures exploiting quantum properties).
The lowest level of this stack consists of quantum hardware, which includes the physical realization of qubits using technologies such as superconducting circuits, trapped ions, quantum dots, or neutral atoms (Bertels et al., 2021).
Above this lies the quantum control level, responsible for precise manipulation and measurement of qubit states. Quantum compilation translates algorithmic descriptions into hardware-specific operations, optimizing quantum circuits to minimize resource usage and mitigate errors. Finally, at the highest level are quantum algorithms and applications, which leverage quantum mechanical principles such as superposition and entanglement to solve complex computational problems efficiently.
This hierarchical model of the quantum computing stack provides a structured perspective for understanding how high-level quantum applications ultimately interact with underlying physical hardware, helping researchers and practitioners navigate the complexities involved in developing quantum computing systems.
@@ 3 Benefits of Quantum Computing
This section discusses key advantages of quantum computing compared to classical computing, emphasizing complexity-theoretic insights, quantum supremacy, and computational speedups. Understanding these benefits motivates ongoing research and development in quantum computing technologies.
3.1 Quantum Complexity Classes and Speedup